Solve the equation
46. 3/4w = -1/2
Any help with this equation would be Greatly appreciated.

Answer by stanbon(75887)

3/4w = -1/2
Multiply both sides by 4/3 to get:
w = (-1/2)(4/3)
w = -2/3
==============
Cheers,
Stan H.